QUESTION 33

  1. Using the given information compute price percentage change if yield increases by 200 basis points.

Yield to Maturity

7%

Macaulay Duration

4.397545

a.

The value of the bond will decrease by 8.22% of its current market value.

b.

The value of the bond will decrease by 8.60% of its current market value.

c.

The value of the bond will increase by 8.60% of its current market value.
Answer #1

Modified duration = macaulay duration/(1+YTM) = 4.397545/(1+0.07)=4.1098

% change in bond price = -modified duration*yield change = -4.1098*0.02=-8.22%

a is correct